There are 8 ounces in a 1/2 pound. How many ounces are in 7 3/4 lbs?
Answer by checkley77(12569)

8/.5=x/7.75
.5x=8*7.75
.5x=62
x=62/.5
x=124 ounces in 7 3/4 pounds.
Proof:
8/.5=124/7.75
.5*124=8*7.75
62=62